OpenStudy (anonymous):

In a dog show, 20% of the dogs are eliminated each day. If the competition starts with 2,000 dogs, to the nearest integer, how many dogs will be present at the start of the 5th day?

OpenStudy (anonymous):

Since 20% = .2, multiply 2000 by .2 five times. Like so, 2000*.2 But do this 5 times, and you minis well say 2000 * (.5)^5

OpenStudy (anonymous):

And then round down

OpenStudy (anonymous):

WOOPS! Take 2000 - (2000*(.2)^5)

OpenStudy (anonymous):

i got 480 as my answer.

OpenStudy (anonymous):

I got 655, but I may have messed up.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

@nickjesus3 i think you may have lol! :)

OpenStudy (anonymous):

oh i read it wrong. just want \(2000\times .8^5\)

OpenStudy (anonymous):

Yeah, I definitely messed up this one... Perhaps 480 is correct

OpenStudy (anonymous):

Its okay i think im right lol but im still not sure !!

OpenStudy (anonymous):

20% decrease every day means 80% remain

OpenStudy (anonymous):

655 looks good to me

OpenStudy (anonymous):

my study guide only shows the answer choices of 259 480 819 1,229

OpenStudy (anonymous):

Yeah, after I realized the equations I put up were wrong I manually did 2000-(2000*.2) = 1600 - (1600*.2) = 1280 - (1280*.2) =....=655

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

hmmmmm.... i wonder whats wrong lol it must be the study guide!! lol or idk what it is

OpenStudy (anonymous):

@hero do you see what the problem is??

hero (hero):

I'm pretty sure @satellite73 is correct.

OpenStudy (anonymous):

yeah so am i

hero (hero):

But I think the question wants \[2000 \times .8^4\]

OpenStudy (anonymous):

ah so it does!!

OpenStudy (anonymous):

is 819 a possibility?

OpenStudy (anonymous):

yes @satellite73